A majorette in a parade is performing some acrobatic twirlings of her baton. Assume that the baton is a uniform rod of mass 0.120 kg and length 80.0 cm .

A majorette in a parade is performing some acrobat
Answer #1

Mass of the rod = 0.12 kg

Moment of inertia of the rod about the centre = I = mL2/12 = (0.12 x(0.8)2)/12 = 0.0064 kg-m2

Angular velocity of the rod = 3 rad/s

Hence, angular momentum of the rod = I\omega = 0.0064 x 3 = 0.0192kg.m2/s